OpenStudy (anonymous):

can anyone help me with the equation 4x-3y=2 ,find y

OpenStudy (anonymous):

please help whoever can show me the equation it simplifys to

OpenStudy (darkigloo):

you want to isolate y. So subtract 4x on both sides.

OpenStudy (darkigloo):

-3y = 2 - 4x you get that right?

OpenStudy (darkigloo):

Divide both sides by -3. y = (2-4x)/-3 y = -2 + 4x/3

OpenStudy (anonymous):

THANKYOU